# LeapFrog: Accelerating Multiscale Materials Simulations with Machine Learning

### v2

@@ Acknowledgements
We acknowledge support from the Mitacs Organization through its Accelerate Fellowship program and Hydro-Québec. We also acknowledge the computational resources made available by the Digital Research Alliance of Canada.
@@
D.P. designed and trained the ML model, generated the data for model training and validation, implemented the LeapFrog algorithm, performed analyses of the results and wrote the paper. M.G. provided PF AMR code, found simulation parameters that demonstrated significant side-branching. M.G. and N.P advised the ML model and LeapFrog algorithm design. N.P. supervised all work, and edited the paper.
@@
Data Availability
The datasets used and/or analyzed during the current study available from the corresponding author on reasonable request.
Code Availability
The underlying code for this study and training/validation datasets are not publicly available, but may be made available to qualified researchers on reasonable request from the corresponding author.

### v3

@@ Acknowledgements
We acknowledge support and funding from the Mitacs Organization through its Accelerate Fellowship program and IREQ-Hydro-Québec. We also acknowledge funding from the Canada Research Chairs program, Grant ID 235492 . Finally, we acknowledge the computational resources made available by the Digital Research Alliance of Canada.
@@
D.P. designed and trained the ML model, generated the data for model training and validation, implemented the LeapFrog algorithm, performed analyses of the results and wrote the paper. M.G. provided PF AMR code, found simulation parameters that demonstrated significant side-branching. M.G. and N.P advised the ML model and LeapFrog algorithm design. N.P. supervised all work, and co-wrote the paper.